THz SPECTROSCOPY OF THE a^{1}Δ_{g} STATE OF O^{18}O and ^{18}O_{2}.

\underline{HARSHAL GUPTA}, SHANSHAN YU, BRIAN J. DROUIN, AND CHARLES E. MILLER, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA 91109; HOLGER S. P. MÜLLER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, Köln, Germany.

Submillimeter-wave rotational spectra of O^{18}O and ^{18}O_{2} in the a^{1}Δ_{g} state have been detected in a new static discharge cell useful for studying rare isotopologues of transient molecules. Rotational transition frequencies of both species have been measured through 1.6 THz, to better than 1 part in 10^{7}, allowing precise determination of molecular constants. The first measurements with our static discharge cell provide a check on the isotopic predictions in the a^{1}Δ_{g} state.